OpenStudy (anonymous):

A football field measures 300 feet in length. It can also be expressed as approximately mc019-1.jpg miles. If there are 5,280 feet in 1 mile, which is the most reasonable value of n?

OpenStudy (anonymous):

@Destinymasha @Abhisar

OpenStudy (anonymous):

We have an option of 4 answers, 2, -2, 3, and -3

OpenStudy (anonymous):

We're thinkin -2

OpenStudy (tkhunny):

Multiply by 1 (one) \(300\;ft\cdot\dfrac{1\;mile}{5280\;ft}\) Reconcile and you'll be done.

OpenStudy (anonymous):

^ That didn't help :/

OpenStudy (anonymous):

mc019-1.jpg= \[5.7*10^n\]

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Did you reconcile? Find the matching units and remove them. Gather all the numbers and multiply or divide them.